# “Did They Consent to That?”: Safer Digital Intimacy via Proactive Protection Against Image-Based Sexual Abuse

### v1

@@ 9.1 Technological Recommendations
In response to these narratives, we conducted a systematic analysis, detailed in Appendix B, of safety feature availability in 8 popular platforms. We found that none of them fully implement all of the features (or features were only available in a special mode/for paid users). We emphasize that wider adoption of features is needed, particularly in popularly-used platforms for protect the majority of sexters who prioritize convenience or defer to their recipient’s choice.
Content-Level Data Control & Presets. Participants wanted variations of existing features for data control that better reflects how they navigate trust and to help separate contexts.
@@
A few participants suggested building a designated app for storing intimate content where content would be stored locally in an encrypted format and that they could then selectively share content from. Although this sounds similar to vault apps, [^8] Vault apps are designated apps for storing sensitive content such as images or documents but may offer varying levels of security [62]. They are often disguised as other apps to avoid discovery when device sharing. only one participant had used one. Similarly, a few others wanted a mechanism for sharing a photo album that they could easily revoke access to: “I have never seen any app…that would allow you to like, you know, sort of create a photo album inside of your conversation” (P7). This suggestion is similar to Grindr’s Album feature [31], which multiple participants wished more platforms would adopt.
@@ 9.2 Research Directions
After being provided a high-level description (see Appendix C.4), participants had mixed reactions of both interest and skepticism. A few expressed enthusiasm about this functionality enabling them to more easily find NDII content and perpetrators but others had reservations around the security of fingerprint storage and whether the fingerprint itself could be used to identify or otherwise violate the image owner’s privacy: “what if it could be traced back to me?” (P10). Meanwhile, a few participants commented that identifying the content alone would not be useful since barriers exist to removing nonconsensually shared content on many platforms.
@@
Broadly, there is a need for visible and accessible resources for people to learn how to protect themselves against NDII. Presently, the most common advice participants received from society, media, and others was not to share any intimate content. This was viewed as impractical, stigmatizing, and responsible for creating a culture of victim blaming because: “everyone does it… it’s a part of how we sort of sexually express ourselves, or make money… If something does happen, and it does go badly, then you feel like, 'Oh, well, everyone’s just saying, I shouldn’t have done it.'So somehow, it’s my fault even though it’s obviously not” (P31).
@@ Acknowledgments
We thank our participants for sharing their experiences with us. This work was supported in part by the European Sex Workers Rights Alliance, the Max Planck Institute for Software Systems, and the U.S. National Science Foundation under award CNS-2206950. We additionally thank Allison McDonald for feedback on an earlier draft of this work.
Appendix A Sharing Technologies and Use Cases
Throughout the interviews, participants mentioned over 40 different platforms for sharing and storing intimate content in the following categories: messaging, social media, adult content platforms, dating/hookup apps, cloud/file share platforms, and email. In Table 1, we elaborate on the use cases people mentioned in relation to each category of sharing method and the number of different platforms that were mentioned per category. This leads us to emphasize that any platform that enables the storing and sharing of photo/video content is likely being used for intimate content.
Appendix B Feature Availability Across Platforms
We conducted an analysis of feature availability in the following popular platforms: Apple iMessage, Discord, Google Messages, Instagram (direct messages), Signal, Snapchat, and WhatsApp. We selected features based on what participants reported using and what they shared that they wanted in a platform for sharing intimate content. This included: screenshot prevention, download prevention, screenshot notification, watermarking, mask on send, automatic metadata removal, blur identifying features (editing assistance with blurring faces, tattoos, backgrounds, etc.), message unsend, expiring messages. The results are displayed in Table B. We designated a platform as providing a feature if it was available in some variant. This includes when it is only available in a specific mode, such as "vanish mode" in Instagram or a hidden/locked chat in Whatsapp. To assess the availability of features, we relied on published documentation and user guides from the platforms themselves. In some cases, we directly downloaded and accessed the platforms we surveyed to confirm the availability/unavailability of a feature when needed.
Sharing Method Usage # of Platforms
Messaging
Sharing in 1-1 conversation or with groups of recipients
9
Social Media
Sharing in 1-1 conversation (DMs), with groups (posting or DMs), for marketing (commercial)
10
Adult content platforms
Sharing content on platforms, livecamming
11
Dating / hookup apps
Finding sexting partners, sharing in-app intimate content, posting semi-nude/lewd photos in profile
10
Cloud/file share platforms
Storing and directly sharing content with recipients
4
Email
Sharing larger files and/or links to files, occasionally used as an alternative to primary sharing technology
N/A
@@ Appendix B Feature Availability Across Platforms
l|c|c|c|c|c|c|c|c|c & Apple iMessage Discord Google Messages Grindr Instagram DM Signal Snapchat WhatsApp

### v2

@@ 9.1 Technological Recommendations
In response to these narratives, we conducted a systematic analysis of 8 popular platforms (see Appendix A) of the availability of safety features participants desired. We found that none of them fully implement all of the features (or features were only available in a special mode/for paid users). We emphasize that wider adoption of features is needed, particularly in popularly used platforms, to protect the majority of sexters who prioritize convenience or defer to their recipient’s choice.
Content-Level Data Control & Presets. Participants wanted variations of existing features for data control that better reflect how they navigate trust and to help separate contexts.
@@
A few participants suggested building a designated app for storing intimate content where content would be stored locally in an encrypted format and that they could then selectively share content from. Although this sounds similar to vault apps, [^10] Vault apps are designated apps for storing sensitive content such as images or documents but may offer varying levels of security [62]. They are often disguised as other apps to avoid discovery when device sharing. See [vault] for a more in-depth discussion on the use of vault apps for intimate content storage. only one participant had used one. Similarly, a few others wanted a mechanism for sharing a photo album that they could easily revoke access to: “I have never seen any app…that would allow you to like, you know, sort of create a photo album inside of your conversation” (P7). This suggestion is similar to Grindr’s Album feature [31], which multiple participants wished more platforms would adopt.
@@ 9.2 Research Directions
After being provided a high-level description (see Appendix B.3), participants had mixed reactions of both interest and skepticism. A few expressed enthusiasm about this functionality enabling them to more easily find NDII content and perpetrators but others had reservations around the security of fingerprint storage and whether the fingerprint itself could be used to identify or otherwise violate the image owner’s privacy: “what if it could be traced back to me?” (P10). Meanwhile, a few participants commented that identifying the content alone would not be useful since barriers exist to removing nonconsensually shared content on many platforms.
@@
Broadly, there is a need for visible and accessible resources for people to learn how to protect themselves against NDII. Presently, the most common advice participants received from society, media, and others was not to share any intimate content. This was viewed as impractical, stigmatizing, and responsible for creating a culture of victim blaming because:
“everyone does it… it’s a part of how we sort of sexually express ourselves, or make money… If something does happen, and it does go badly, then you feel like, 'Oh, well, everyone’s just saying, I shouldn’t have done it.'So somehow, it’s my fault even though it’s obviously not” (P31).
@@ Acknowledgments
We thank our participants for sharing their experiences with us. This work was supported in part by the European Sex Workers Rights Alliance, the Max Planck Institute for Software Systems, and the U.S. National Science Foundation under award CNS-2206950. Lucy Qin was additionally supported by the NSF Graduate Research Fellowship. We thank Allison McDonald for feedback on an earlier draft of this work.
Appendix A Feature Availability Across Platforms
We conducted an analysis of feature availability in 8 popular platforms in November 2023. We selected features based on what participants reported using and what they shared that they wanted in a platform for sharing intimate content. The results are displayed in Table 2. We designated a platform as providing a feature if it was available in some variant. This includes when it is only available in a specific mode, such as "vanish mode" in Instagram or a hidden/locked chat in Whatsapp. To assess the availability of features, we relied on published documentation and user guides from the platforms themselves. In some cases, we directly downloaded and accessed the platforms we surveyed to confirm the availability/unavailability of a feature when needed.
@@ Appendix A Feature Availability Across Platforms
Apple iMessage
Discord
Google Messages
Grindr
Instagram DM
Signal
Snapchat
WhatsApp
@@
Appendix B Methods (Continued)
B.1 Recruitment Graphics
The following recruitment graphics were used to recruit recreational participants (Figure 1) and commercial participants (Figure 2).
B.2 Privacy and Anonymity.
To protect the privacy and anonymity of the participants, we used end-to-end encrypted platforms for the interviews, provided advice on how to sign up with an encrypted (and anonymous) email address if preferred, and used Calendly to schedule interviews, which did not require any legal information or login. The payment options could be anonymously received as well. Interviews were end-to-end encrypted and conducted either using Webex or Zoom. Participants were given the option of conducting the interview as a video call, audio call, or text chat. Sex-working participants were given the option of being interviewed by a current or former sex worker.
Figure 1: Recruitment graphic for recreational sexters
Figure 2: Recruitment graphic for sex workers
B.3 Interview Protocols
Our interview protocols are available at the following link along with our codebook:
bit.ly/Tech-IBSA-Research-Protocols
The audio and video interviews lasted an average of 47 minutes. During the interviews, we explicitly asked participants about the usefulness of a selected set of features based on prior work and existing media advice guides. These features included: end-to-end encryption, expiring messages, screenshot notifications, digital fingerprinting, watermarking, photo editing tools, and metadata removal. We wrote a high-level explanation of each feature that interviewers read to participants who had not previously heard of them or were only vaguely familiar with them. A document containing these explanations are included in the same repository.